# QUITE: A Query Rewrite System Beyond Rules with LLM Agents

### v2

@@ 6. Query Hint Recommender
Motivation. Query rewrite applies high-level transformations to give the DBMS optimizer a better starting point for plan generation. While this improved starting point is beneficial, the optimizer, which dictates the finer-grained execution operations, can still generate suboptimal plans. This sub-optimality arises because the optimizer’s decisions rely on cost models that are often inaccurate due to outdated statistics and simplifying assumptions (e.g., statistical independence of columns) (Stillger et al., 2001; Marcus et al., 2019; Chen et al., 2023). To exert more fine-grained control over plan selection, we propose embedding query hints (Marcus et al., 2021) within rewritten queries. In this work, we introduce a dedicated hint base and use LLMs to adaptively select and incorporate effective hints into rewritten queries.
@@ 6.1. Hint Base Construction
Most modern database systems (e.g., PostgreSQL, MySQL, Oracle) support query hints, either inherently or by extensions (NTT OSS Center DBMS Development and Support Team, 2012; Oracle Corporation, 2024, 2019). In this work, we use PostgreSQL’s pg_hint_plan extension (NTT OSS Center DBMS Development and Support Team, 2012) to implement our query hint system. Our system supports a broader range of hints (e.g., disabling materialization and setting specific cardinality values), whereas existing works (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023; Negi et al., 2021) are limited to Boolean hints.
@@
Introducing a Novel CTE Hint. To provide more precise control over Common Table Expression (CTE) execution and address associated performance issues, we introduce a new NO_MATERIALIZE hint. By default, PostgreSQL materializes CTEs (pos, [n.d.]a). This approach can lead to significant performance degradation, particularly when CTEs operate on small datasets or are executed infrequently (El-Helw et al., 2015; Zhou et al., 2007), due to the overhead of writing and reading intermediate results. Furthermore, the decision to materialize a CTE, as opposed to inlining it, directly impacts the optimizer’s ability to perform effective predicate pushdown into subqueries, especially after query rewrite. Our NO_MATERIALIZE hint directs the optimizer to inline the targeted CTE, expanding its definition into the main query as if it were a FROM-clause subquery. This hint enhances query performance by eliminating materialization overhead and enabling more effective plan optimizations, such as improved predicate pushdown. Ultimately, these selected and newly introduced hints constitute our query hint base $\mathcal{H}$, detailed in Table 2.
@@ 6.2. Hint Injection
This section details our approach to using LLMs for identifying operations that could benefit from hints and then recommending appropriate ones. Unlike existing methods (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023; Negi et al., 2021) that recommend hints at the query level, where a single hint applies to all operations within that query, our strategy targets a more granular level. We found that query-level hints can be suboptimal, as a hint beneficial for one operation might negatively impact others. Therefore, we apply hints at a finer-grained level, such as to individual tables or specific operations within a query, to maximize positive impacts.
Building on the validated rewrite of $Q^{\prime}$, we now turn to refining its physical execution plan through targeted hint injection. First, given the physical plan of $Q^{\prime}$, database semantics and data distribution statistics, we identify potential improvements that can be achieved through hints: (1) For every cardinality estimation in the plan, an LLM judges its reasonableness. If an estimation is deemed unreasonable, the LLM provides a recommended range and justification. (2) For all join operations, the LLM assesses whether the selected join operator (such as hash, merge, or nested-loop join) is suitable given the available statistics. Second, an LLM is then invoked to construct the most effective hints based on the obtained analysis and the query plan. It recommends join methods for specific tables and rectifies cardinality estimations. Regarding NO_MATERIALIZE, we apply this hint only if a CTE appears exactly once or processes a small, simple dataset. Finally, the proposed hints are then injected into $Q^{\prime}$, producing the final optimized query $Q^{*}$.
Our LLM-guided hint injection, unlike general existing methods, pinpoints costly operations and generates finer-grained hint recommendations. This enhances the explainability of the recommendations and leads to further improvement. Moreover, in contrast to ML-based hint recommendation approaches (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023; Negi et al., 2021) that rely on pre-trained models and multi-plan enumeration, our method requires zero training yet surpasses their performance.
@@ 7.1. Experimental Set up
Testbed. Our experiments were conducted on a server featuring a 32-core Intel Xeon Platinum 8352V CPU, 251 GB of RAM, a 942 GB SSD, and PostgreSQL v14.13. We utilize a NVIDIA A100-SXM4 GPU with 40 GB of memory to run the rewrite process of LLM-R2.
Workloads. We employ three widely recognized datasets for our evaluations. (1) TPC-H (TPC, nd) is a well-known OLAP benchmark containing 62 columns and 22 query templates. We generate 63 queries from these templates, excluding Q15 due to its use of CREATE VIEW, a format unsupported by the rewrite engines of two baseline methods (Zhou et al., 2021; Li et al., 2024). We used a scale factor (SF) of 10. (2) DSB (Ding et al., 2021) is adapted from the TPC-DS benchmark (Nambiar and Poess, 2006) and is notable for its complex data distribution and challenging long-context query templates, which consists of 52 query templates. We generate 156 queries based on these templates with $SF=10$. (3) Calcite (apa, 2021) is a real-world workload used to evaluate rewrite rules. We randomly select 58 queries and generate 10G of data with uniform distributions following previous work (Dong et al., 2023).
@@
  • LearnedRewrite (LR) (SIGMOD 2022)  (Zhou et al., 2021) uses a Monte Carlo Tree Search (MCTS) algorithm with a learned cost estimation model to explore a SQL policy tree of rewrite rule orders. We adopt the original implementation and use the cost estimation model provided in their official code repository.
  • LLM-R2 (VLDB 2025) (Li et al., 2024) is a rule-based query rewrite system built on Apache Calcite (Begoli et al., 2018). It uses LLM’s In-Context Learning (ICL) ability to select query rewrite rules according to high quality demonstration queries.
  • R-bot (Sun et al., 2024) is an LLM-based system that leverages an embedded retrieval-augmented generation (RAG) knowledge base and reflective reasoning to select promising rewrite rules.
@@
Table 3. Different Methods on Query Latency (s)
Methods TPC-H (SF=10) DSB (SF=10) Calcite (SF=10)
Mean Median 75th 95th Mean Median 75th 95th Mean Median 75th 95th
Original 69.84 9.64 32.75 300.00 32.62 4.85 10.29 300.00 23.88 2.58 8.64 122.36
LearnedRewrite 37.57 9.97 30.55 202.31 31.93 5.14 15.18 251.42 22.88 2.32 8.09 122.58
LLM-R2(GPT-4o) 57.09 9.30 30.50 300.00 10.53 4.11 8.36 32.88 16.98 2.11 7.75 50.97
LLM-R2(Claude-3.7) 57.68 9.89 30.07 300.00 9.11 3.82 8.24 27.48 16.96 2.55 8.19 50.07
LLM-R2(DS-R1) 56.73 9.25 29.33 300.00 9.16 3.84 8.21 24.21 16.90 2.05 7.71 49.91
R-Bot (GPT-4o) 32.64 9.48 26.25 172.20 21.75 4.26 8.80 119.19 22.21 2.34 8.35 122.14
R-Bot (Claude-3.7) 33.89 9.71 29.32 195.51 19.89 4.27 8.99 112.24 23.38 2.60 8.64 122.95
R-Bot (DS-R1) 33.70 9.88 28.38 176.90 21.21 4.44 9.11 113.91 22.71 2.52 8.43 122.89
LLM Agent (GPT-4o) 69.03 10.00 27.02 300.00 33.44 4.39 9.65 300.00 15.47 2.03 7.32 35.34
LLM Agent (DS-R1) 42.79 9.94 27.95 300.00 13.53 3.87 7.93 30.18 15.18 1.65 7.55 36.70
LLM Agent (DS-V3) 60.12 9.22 29.92 300.00 32.34 4.19 9.99 300.00 16.49 2.53 7.72 45.15
LLM Agent (Claude-3.7) 57.46 11.04 36.99 300.00 25.32 4.38 8.84 300.00 15.67 1.59 7.68 44.54
@@
(a) TPC-H Benchmark
(b) DSB Benchmark
(c) Calcite Benchmark
@@
QUITE Settings. We use DeepSeek-R1 as the reasoning agent due to its strong analysis capability, and Claude-3.7-Sonnet as the rewrite, decision and assistant agents as default. A temperature of 0 is applied to all LLMs to ensure output stability  (DeepSeek-AI, 2025; Anthropic, 2025). The number of iterations is limited to 2 to prevent excessive FSM rewrite cycles. All LLM-involved experiments employ the identical temperature configurations used in our approach.
@@
Query Evaluation Approach. Before each query, we restart the evaluation database to clear all caches. The query execution process involves an initial warm-up run (to mitigate cold start impacts), followed by three measured executions. We report the average execution time of these three runs. Queries that run for more than 300 seconds are terminated and their time is recorded as 300 seconds (Li et al., 2024), to reflect impractical execution durations. If a rewritten query produces different output than the original or contains syntax errors, its equivalence metric is marked as False. To simulate a realistic query optimization scenario when dealing with such nonequivalent queries, we assign the rewritten query the execution time of its original counterpart.
@@ 7.2. Performance Comparison
Query Execution Latency. As presented in Table 3, queries optimized by QUITE demonstrate significantly reduced execution times, consistently outperforming all baseline methods. Across the TPC-H, DSB, and Calcite benchmarks, QUITE achieves substantial average execution time reductions. Specifically, QUITE reduces execution time by 31.9%, 81.7%, and 56.4% compared to LR; 54.9%, 35.8%, and 41.0% over LLM-R2; and 40.2%, 56.8%, and 34.3% against LLM Agent. When evaluated against R-Bot, QUITE achieves reductions of 21.6%, 70.6%, and 55.1%. QUITE’s advantages are further illustrated by the rewritten improvement rates shown in Figure 6, where QUITE achieves the highest ratios of 31.7%, 37.8%, and 34.5% on TPC-H, DSB, and Calcite, respectively, indicating strong generalization capabilities across different workloads.
The optimization benefits of QUITE over baselines are more pronounced on complex datasets. Notably, the improvement margin of QUITE over LR is substantially larger for the DSB dataset compared to the TPC-H and Calcite datasets. This is primarily because: (1) The LR we use is trained on the TPC-H, LR can proficiently identify effective rewrite rules and navigate their constrained search spaces. This existing proficiency limits the remaining optimization potential, thereby reducing the margin for QUITE to demonstrate substantial further improvements. (2) LR employs a greedy search within its MCTS algorithm, which proves less effective in the expansive search spaces of complex datasets like DSB, unlike its performance on simpler datasets (e.g., Calcite) with smaller search spaces and inherently less optimization potential.
@@
Rewritten Equivalence. Figure 6 details the number of equivalent and improved query rewrites produced by each method across the datasets. Notably, direct rewrites from the LLM Agent exhibit significant semantic equivalence issues. In contrast, QUITE achieves substantially higher equivalence rates–attaining 100%, 96.8%, and 98.3% on TPC-H, DSB, and Calcite, respectively. This superior performance is attributable to the multi-stage correction mechanisms integrated into QUITE’s iterative FSM-based process.
@@
7.3. Ablation Study
Rewrite Middleware. To assess the hybrid SQL corrector, we removed it and evaluated its impact on query equivalence and execution efficiency using the DSB dataset. As shown in Figure 8, judge accuracy measures the proportion of correct executions among all decisions made by the tool or LLM while coverage ratio indicates the proportion of queries for which the tool or LLM was able to make a definitive decision.Its removal significantly degrades performance: nonequivalent queries increase from 5 to 11, unchanged queries by 13, and improved queries drop by 17. These results highlight the corrector’s essential role in ensuring query quality and efficiency.
We also evaluate the impact of the structured knowledge component. As shown in Table 4, QUITE consistently outperforms its variant without this module across all metrics. The improvement stems from high-quality Q&A pairs and domain knowledge embedded in the knowledge base, which effectively guide more optimized rewrites.
Table 4. Ablation Study of Structured Knowledge Base
Methods DSB (SF=10)
Mean Median 75th 95th
Without Knowledge 9.96 3.94 8.49 24.81
With Knowledge 6.08 3.78 7.98 21.43
@@ 7.3. Ablation Study
LLM-Agent-based Query Rewrite FSM. To evaluate the impact of the FSM framework within our system, we performed two modifications: (1) the loop responsible for equivalence checking was removed, and (2) the decision stage was eliminated. These changes transformed the process into a linear, end-to-end pipeline while retaining all components and agent tools involved in query rewriting. At the same time, to further assess the impact of the MDP-based decision-making on the Reasoning agent’s capabilities, we conducted an ablation study by removing the MDP component. As shown in Table 5, QUITE outperforms its variant across all evaluation metrics. Notably, the average execution time with the MDP is reduced by 52.2% compared to the version without it, demonstrating that the MDP significantly enhances the effectiveness of the Reasoning agent.
Table 6. Ablation Study of QUITE using Different LLMs
Methods DSB (SF=10)
Mean Median 75th 95th
Deepseek-R1 + GPT-4o 31.38 4.64 9.56 300.00
Deepseek-R1 + Deepseek-V3 28.07 4.43 9.07 300.00
Deepseek-R1 + Claude-3.7 6.08 3.78 7.98 21.43
Table 7. Ablation Study of Hint Construction Methods
Methods DSB (SF=10)
Mean Median 75th 95th
QUITE (without hints) 6.08 3.78 7.98 21.43
QUITE (with Bao’s Hint Base) 6.93 4.65 8.46 20.41
QUITE (with Our Hint Base) 5.85 3.35 7.63 19.97

### v3

@@ 6. Query Hint Recommender
Query rewrite applies high-level transformations to give the DBMS optimizer a better starting point for plan generation. While this improved starting point is beneficial, the optimizer, which dictates the finer-grained execution operations, can still generate suboptimal plans. This sub-optimality arises because the optimizer’s decisions rely on cost models that are often inaccurate due to outdated statistics and simplifying assumptions (e.g., statistical independence of columns) (leo; neo; base). To exert more fine-grained control over plan selection, we propose embedding query hints (bao) within rewritten queries. In this work, we introduce a dedicated hint base and use LLMs to adaptively select and incorporate effective hints into rewritten queries.
@@ 6.1. Hint Base Construction
Most modern database systems (e.g., PostgreSQL, MySQL, Oracle) support query hints, either inherently or by extensions (pg_hint_plan; mysql_hints; oracle_hints). In this work, we use PostgreSQL’s pg_hint_plan extension (pg_hint_plan) to implement our query hint system. Our system supports a broader range of hints (e.g., disabling materialization and setting specific cardinality values), whereas existing works (bao; autosteer; fastgres; steeringqo) are limited to Boolean hints.
@@
Introducing a Novel CTE Hint. To provide more precise control over Common Table Expression (CTE) execution and associated performance issues, we introduce a new NOT_MATERIALIZE hint. By default, PostgreSQL materializes CTEs (postgresqlcte), which can lead to significant performance degradation, particularly when CTEs operate on small datasets or are executed infrequently (cte1; cte2), due to the overhead of writing and reading intermediate results. Furthermore, the decision to materialize a CTE, as opposed to inlining it, directly impacts the optimizer’s ability to perform effective predicate pushdown into subqueries, especially after query rewrite. Our NOT_MATERIALIZE hint directs the optimizer to inline the targeted CTE, expanding its definition into the main query as if it were a FROM-clause subquery. This hint enhances query performance by eliminating materialization overhead and enabling more effective plan optimizations, such as improved predicate pushdown. Ultimately, these selected and newly introduced hints constitute our query hint base $\mathcal{H}$, detailed in Table 3.
@@ 6.2. Hint Injection
This section explains how we use LLMs to identify operations that could benefit from hints and recommend appropriate ones. Prior query-level methods (bao; autosteer; fastgres; steeringqo) apply a single hint to the entire query, but we observe that such a hint can be suboptimal, as a hint that benefits one operation may degrade others. To overcome this limitation, our strategy instead applies hints at the operator level, such as to individual tables or specific operations within a query, to maximize their positive effects.
Building on the validated rewrite of $Q^{\prime}$, we now turn to refining its physical execution plan through targeted hint injection. First, given the physical plan of $Q^{\prime}$, database semantics and data distribution statistics, we identify potential improvements that can be achieved through hints: (1) For every cardinality estimation in the plan, an LLM judges its reasonableness. If an estimation is deemed unreasonable, the LLM provides a recommended range and justification. (2) For all join operations, the LLM assesses whether the selected join operator (such as hash, merge, or nested-loop join) is suitable given the available statistics. Second, the Assistant Agent processes these justifications, verifies their consistency against plan statistics, and constructs the final set of hints before forwarding them to the injection function. Regarding NOT_MATERIALIZE, we apply this hint only if a CTE appears exactly once or processes a small, simple dataset. Finally, the proposed hints are then injected into $Q^{\prime}$, producing the final optimized query $Q^{*}$.
Unlike general existing methods (autosteer; bao; steeringqo; fastgres), our LLM-guided hint injection pinpoints costly operations and generates finer-grained hint recommendations. This enhances the explainability of the recommendations and leads to further improvement. Moreover, in contrast to ML-based hint approaches (autosteer; bao; steeringqo; fastgres) that rely on pre-trained models and multi-plan enumeration, our method requires zero training yet shows a competitive performance within a small subset of query hints.
@@ 7. Experiment
In this section, we conduct a comprehensive evaluation of QUITE through performance comparisons, ablation studies, robustness assessments, analyses of rewrite behaviors and cost evaluations.
@@ 7.1. Experimental Set up
Testbed. Our experiments were conducted on a server featuring a 32-core Intel Xeon Platinum 8352V CPU, 251 GB of RAM, a 942 GB SSD, and PostgreSQL v14.13.
Workloads. We evaluate QUITE on standard OLAP benchmarks with long-running analytical queries, aligning with our deployment scenarios. We employ three widely recognized datasets and one LLM-synthesized dataset to evaluate all of the approaches. (1) TPC-H (TPC_H_toolkit) is a well-known OLAP benchmark containing 62 columns and 22 query templates. We generate 63 queries from these templates, excluding Q15 due to its use of CREATE VIEW, a format unsupported by the rewrite engines of two baseline methods (lr; llm-r2). We used a scale factor (SF) of 10. (2) DSB (ding2021dsb) is adapted from the TPC-DS benchmark (tpcds) and is notable for its complex data distribution and challenging long-context query templates, which consists of 52 query templates. We generate 156 queries based on these templates with $SF=10$. (3) Calcite (apache_calcite_test) is a real-world workload used to evaluate rewrite rules in Apache Calcite (apache). We randomly select 58 queries and generate 10G of data with uniform distributions following previous work (slabcity). (4) StackOverflow (StackOverflow-Math) (stackoverflow-math) is a 13.8 GB dataset from the Mathematics community on the Stack Exchange network [^1] Stack Overflow is the flagship site of the network (sqlstorm), containing math-related posts and metadata as of October 3, 2024. Following  (sqlstorm), we generate queries using predefined complexity prompts (P1–P5) and obtain 43 valid queries after dialect correction.
@@
  • LearnedRewrite (LR) (SIGMOD 2022)  (lr) uses a Monte Carlo Tree Search (MCTS) algorithm with a learned cost estimation model to explore a SQL policy tree of rewrite rule orders. We adopt the original implementation and use the cost estimation model provided in their official code repository.
  • LLM-R2 (VLDB 2025) (llm-r2) is a rule-based query rewrite system built on Apache Calcite (apache). It uses LLM’s In-Context Learning (ICL) ability to select query rewrite rules according to high quality demonstration queries.
  • R-bot (VLDB 2025) (r-bot) is an LLM-based system that leverages an embedded retrieval-augmented generation (RAG) knowledge base and reflective reasoning to select promising rewrite rules.
@@
Table 4. Query Latency of Different Methods (s)
Methods TPC-H (SF=10) DSB (SF=10) Calcite (SF=10) StackOverflow (SF=10)
Mean Median 75th 95th Mean Median 75th 95th Mean Median 75th 95th Mean Median 75th 95th
Original 69.84 9.64 32.75 300.00 32.62 4.85 10.29 300.00 23.88 2.58 8.64 122.36 48.02 9.47 20.80 300.00
LearnedRewrite 37.57 9.97 30.55 202.31 31.93 5.14 15.18 251.42 22.88 2.32 8.09 122.58 46.25 9.43 18.35 300.00
LLM-R2(GPT-4o) 57.09 9.30 30.50 300.00 10.53 4.11 8.36 32.88 16.98 2.11 7.75 50.97 47.08 9.19 16.06 300.00
LLM-R2(Claude-3.7) 57.68 9.89 30.07 300.00 9.11 3.82 8.24 27.48 17.80 2.30 7.83 57.28 46.97 8.40 16.30 300.00
LLM-R2(DS-R1) 56.73 9.25 29.33 300.00 9.16 3.84 8.21 24.21 16.90 2.05 7.71 49.91 43.42 9.21 14.14 300.00
R-Bot (GPT-4o) 32.64 9.48 26.25 172.20 21.75 4.26 8.80 119.19 22.21 2.34 8.35 122.14 46.43 9.37 16.02 300.00
R-Bot (Claude-3.7) 33.89 9.71 29.32 195.51 19.89 4.27 8.99 112.24 23.38 2.60 8.64 122.95 52.06 9.52 19.25 300.00
R-Bot (DS-R1) 33.70 9.88 28.38 176.90 21.21 4.44 9.11 113.91 22.71 2.52 8.43 122.89 50.57 9.46 17.39 300.00
LLM Agent (GPT-4o) 69.03 10.00 27.02 300.00 33.44 4.39 9.65 300.00 15.47 2.03 7.32 35.34 50.90 7.96 26.73 300.00
LLM Agent (DS-R1) 42.79 9.94 27.95 300.00 13.53 3.87 7.93 30.18 15.18 1.65 7.55 36.70 27.14 6.38 9.40 272.55
LLM Agent (DS-V3) 60.12 9.22 29.92 300.00 32.34 4.19 9.99 300.00 16.49 2.53 7.72 45.15 54.92 10.33 23.60 300.00
LLM Agent (Claude-3.7) 57.46 11.04 36.99 300.00 25.32 4.38 8.84 300.00 15.67 1.59 7.68 44.54 36.23 7.58 21.86 283.42
@@
(a) TPC-H
(b) DSB
(c) Calcite
(d) StackOverflow
@@
Figure 7. Comparison of Overall Latency (h)
Comparison with GenRewrite (genrewrite). GenRewrite explores query rewriting with natural-language rewrite rules and iterative correction, but the lack of released code prevents direct comparison. At a high level, it accumulates natural-language rewrite knowledge to guide subsequent LLM-based rewriting. In contrast, QUITE formulates query rewrite as a decision process and directly optimizes execution through iterative refinement with database feedback, enabling more robust and data-aware optimizations.
QUITE Settings. We use DeepSeek-R1 as the reasoning agent due to its strong analysis capability, and Claude-3.7-Sonnet as the rewrite, decision, and assistant agents by default. A temperature of 0 is applied to all LLMs to ensure output stability (DeepSeek_doc; anthropic_doc). The number of iterations is limited to 2 to prevent excessive FSM rewrite cycles. All LLM-involved experiments employ the identical temperature configurations used in our approach.
@@
Query Evaluation Approach. Before each query, the database is restarted to clear caches. The query execution process involves an initial warm-up run to mitigate cold start impacts, followed by three measured executions. We report the average execution time of these three runs. Queries exceeding 300 seconds are terminated and recorded (llm-r2; r-bot), with capped values included in all latency statistics (mean, median, 75th, and 95th percentiles). Rewritten queries with syntax errors or output mismatches are marked as non-equivalent. In such cases, the execution time of the original query is used. If a rewrite times out, we recheck equivalence on a smaller scale: non-equivalent rewrites fall back to the original, while equivalent rewrites are executed with capped latency.
@@ 7.2. Performance Comparison
Query Execution Latency. As shown in Table 4, queries optimized by QUITE demonstrate significantly reduced execution times, consistently outperforming all baseline methods. Across the TPC-H, DSB, Calcite and StackOverflow benchmarks, QUITE achieves substantial average execution time reductions. Specifically, QUITE reduces execution time by 31.9%, 81.7%, 56.4% and 72.7% compared to LR; 54.9%, 35.8%, 41.0% and 71.0% over LLM-R2; and 40.2%, 56.8%, 34.3% and 53.5% against LLM Agent. When evaluated against R-Bot, QUITE achieves reductions of 21.6%, 70.6%, 55.1% and and 72.8%. QUITE’s advantages are further illustrated by the rewritten improvement rates shown in Figure 6, where QUITE achieves the highest ratios of 31.7%, 37.8%, 34.5% and 65% on TPC-H, DSB, and Calcite, respectively, indicating strong generalization capabilities across different workloads.
The optimization benefits of QUITE are more pronounced on complex datasets. Notably, on the LLM-generated StackOverflow workload, we did not curate queries to disadvantage baselines. Since many LLM-generated queries use broad SQL dialects that are not directly supported by the strict syntax of Apache Calcite used in rule-based pipelines (lr; llm-r2; r-bot), we minimally normalized only such dialect mismatches (e.g., LISTAGG vs. STRING_AGG), discarding unresolvable cases. Under this controlled setup, rule-based pipelines remain bottlenecked by dialect constraints and limited rule and pattern coverage on complex structures, while QUITE is more robust and delivers larger latency gains.
Table 5. Unified Ablation Study Results on DSB (SF=10)
Metrics Structured Knowledge Base Different LLMs Hint Construction Methods
w/o Q&A Units Raw Q&A Units Filtered Q&A Units DS-R1 + GPT-4o DS-R1 + DS-V3 DS-R1 + Claude-3.7 DS-V3 + Claude-3.7 Claude-3.7 + Claude-3.7 w/o Hints Bao’s Hint Base Full Hints (no GUC) Ours + Scan Our Hint Base
Mean 9.96 9.10 6.08 31.38 28.07 6.08 23.41 12.76 6.08 6.93 7.06 6.04 5.85
Median 3.94 3.85 3.78 4.64 4.43 3.78 4.32 4.01 3.78 4.65 4.02 3.89 3.35
75th 8.49 8.10 7.98 9.56 9.07 7.98 9.64 8.74 7.98 8.46 8.64 8.06 7.63
95th 24.81 23.01 21.43 300.00 300.00 21.43 148.47 58.38 21.43 20.41 22.52 20.29 19.97
@@
Overall Latency. As shown in Figure 7, we evaluate overall execution time on TPC-H (SF=50) and DSB (SF=50) with a per-query time limit of 8000 seconds, using Claude-3.7-Sonnet across all baselines. On TPC-H and DSB, QUITE reduced execution time by 71.4% and 53.6% compared with the original queries, outperforming LR (47.6%, 53.4%), LLM-R2 (71.9%, 20.0%), and R-Bot (29.1%, 45.5%), respectively. These results demonstrate QUITE’s strong end-to-end advantage. QUITE’s rewrite time on the DSB dataset is relatively longer, which can be attributed to two main factors: first, DeepSeek-R1 requires more time to process difficult problems; second, for complex queries, the decision agent often performs an additional iteration to explore further optimizations.
Rewritten Equivalence. Figure 6 details the number of equivalent and improved rewrites produced by each method across the datasets. Notably, direct LLM-agent rewrites, especially by Claude-3.7-Sonnet on StackOverflow, often violate equivalence due to subtle semantic drift during query restructuring. In contrast, QUITE achieves substantially higher equivalence rates, attaining 100%, 96.8%, 98.3% and 90.7% on TPC-H, DSB, Calcite and StackOverflow. In addition, rule-based methods suffer from incomplete semantic equivalence, as many rules in their rewrite engines (e.g., Apache Calcite (apache)) lack formal verification due to the inherent difficulty of equivalence checking (wetune; sqlsolver). Despite this, our approach achieves the highest equivalence rate across all benchmarks, advancing the state of the art.
7.3. Ablation Study
Rewrite Middleware. First, we ablated the hybrid SQL corrector to measure its effect on equivalence and execution. Figure 9 reports Judge Accuracy (the fraction of correct decisions) and Coverage Ratio (the fraction of queries with a definitive decision). Its removal significantly degrades performance: nonequivalent queries increase from 5 to 11, unchanged queries increase by 13, and improved queries drop by 17. These results highlight the corrector’s essential role in ensuring query quality and efficiency.
We also evaluate the impact of the structured knowledge base. As shown in Table 5, QUITE consistently outperforms its variant without this module across all metrics. The improvement stems from high-quality domain knowledge embedded in the knowledge base, which effectively guide more optimized rewrites. To justify the LLM-based filtering, we compare the 241 filtered units against all 3,432 collected units. The filtered version delivers significantly stronger execution performance: five queries achieve a $2\times$$5\times$ speedup, and more than eight queries exceed $5\times$. This confirms that consensus-aligned, high-quality knowledge provides significantly more effective guidance for LLM rewriting, enabling stable and efficient query optimization.